NEED HELP!!!!

American Food Services, Inc., acquired a packaging machine from Barton and Barton Corporation. Barton and Barton completed construction of the machine on January 1, 2021. In payment for the $4 million machine, American Food Services issued a four-year installment note to be paid in four equal payments at the end of each year. The payments include interest at the rate of 10%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Prepare the journal entry for American Food Services’ purchase of the machine on January 1, 2021.
2. Prepare an amortization schedule for the four-year term of the installment note.
3. Prepare the journal entry for the first installment payment on December 31, 2021.
4. Prepare the journal entry for the third installment payment on December 31, 2023.
Answer #1
Date Account titles and explanation Debit Credit
1 01-Jan-21 Equipment (packing machine) $4,000,000
   Long term note payable $4,000,000
(equipment purchase for note)
2 Present value annuity factor for 10% for 4 years = 3.16987
Note amount $4,000,000
Annuity value $1,261,881 (4,000,000/3.16987)
A B = (A*10%) C D = (C-B) E = (A-D)
Date Opening value of Note Interest Installment paid Reduction in value of note Ending value of Note
31-Dec-21 $4,000,000 $400,000 $1,261,881 $861,881 $3,138,119
31-Dec-22 $3,138,119 $313,812 $1,261,881 $948,069 $2,190,050
31-Dec-23 $2,190,050 $219,005 $1,261,881 $1,042,876 $1,147,174
31-Dec-24 $1,147,174 $114,707 $1,261,881 $1,147,174 $0
3 31-Dec-21 Interest expense $400,000
Long term note payable $861,881
   Cash $1,261,881
(first installment paid)
31-Dec-23 Interest expense $219,005
Long term note payable $1,042,876
   Cash $1,261,881
(third installment paid)

*feel free to ask any doubt related to this question in the comment section. Please note that I've rounded off present value table values to 5 decimal places in absence of any clear instructions. Please change it accordingly if you have different rounding off criteria.

please give your valuable feedback.